Mr Benard is struggling to repay his loan of $ 400,000 with payments of $ 7,200 made monthly in arrears for 6 years.

(a) Calculate the FLAT RATE OF INTEREST per annum

(b) Hence, or otherwise, calculate the APR of Mr Benard's loan

After exactly one , loan company offers to " help" Mr Benard by restructuring his loan with new monthly payments of $ 4,000 made in arrears

(c) Assuming the company charges the same APR as Mr Benard's original loan. Calculate the term of the new loan

( d) Calculate how much more interest in total Mr Benard will pay on his restructured loan than on his original loan
Answer #1

(a) Flat Rate= 4.9333% as calculated below:

B C D E Values 400,000 P P 3 Loan amount 4 Term (years) 5 Therefore, monthly repayment of principal p 6 Monthly payment total

(b) APR is arrived at using the RATE function of Excel at 8.948524% as follows:

foc =RATE(C5, C6,03*-1,0) *12 A D Cell ref Given Given Values 400,000 2 3 Loan amount 4 Term (years) 5 Number instalments 6 M

(c) It is assumed that the question refers to position 'after exactly one year'

The term of new loan is arrived at 177.88 months using the NPER function of Excel

C16 • Foc =NPER(C7/12,015,C14*-1,0,0) B 1 Cell ref Given Given Values 400,000 72 Given RATE 7,200 8.94852443% 2 3 Loan amount

(d): Total payment of new loan = Monthly payment* new term= $4,000*177.881966 = $711,527.86

Total payment for remaining period as per old plan= Monthly payment* Remaining period =$7,200*5*12 = $432,000

Increase in total payment (which is towards interest, hence increase in interest)= $711,567.86-$432,000 = $279,527.86